OpenStudy (ilikestudying140898):

Please help... been stuck on this question forever!! Solve 9x + 5y + 3z = 4 6x + 2y + 4z = 2 -3x + 3y – 7z = 2

OpenStudy (ilikestudying140898):

(-1, 2, 1) (1, -1, 0) (4, 2, 2) (1, 2, -5)

OpenStudy (ilikestudying140898):

and please explain how you did it....so i can do another questions myself :)

OpenStudy (welshfella):

we can do this by the elimination method First multiply the 3rd equation by 3 and add it to the first equation so we get -9x + 9y - 21z = 6 ( 3rd one * 3) 9x + 5y + 3z = 4 ADD: 14y - 18z = 1o we have eliminated the term in x understand so far?

OpenStudy (ilikestudying140898):

omg!!!! thank you thank you!!!! helped sooo faar!!

OpenStudy (welshfella):

yw Now the next step is to get another equation in x and y only We can do this by multiplying the 3rd equation by 2 and then adding it the the second equation can you try doing that?

OpenStudy (ilikestudying140898):

@welshfella so i the answer (4,2,2)... i guess...omg i am lost...i feel like a dumb person :( gosh i am really bad at systems of equations :<